OpenStudy (anonymous):

Find the solution of the given differential equation satisfying the indicated initial condition. y’ = (ln^2)y, y(2) = 4

terenzreignz (terenzreignz):

Can't guarantee i can do this, but is it \[\Large y\prime = \ln^2(y)\]

OpenStudy (anonymous):

yess my bad

terenzreignz (terenzreignz):

How do you usually do this, with series?

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

\[\frac{dy}{dx}=\ln^2y\] Now, do you mean \(\ln^2y=\ln(\ln y)\), or \(\ln^2y=(\ln y)^2\)? There's some ambiguity with the notation.

OpenStudy (anonymous):

they way its written is (ln2)y

OpenStudy (anonymous):

Well that explains it. \(\ln2\) is a coefficient, approximately 0.69315. The differential equation is separable: \[\frac{dy}{dx}=\ln2 ~y\\ \frac{1}{y}~dy=\ln2~dx\] Integrate both sides: \[\int\frac{1}{y}~dy=\int\ln2~dx\\ \ln|y|=\ln2~x+C\\ y=e^{\ln2~x+C}\\ y=Ce^{\ln 2 ~x}\\ y=C\left(e^{\ln 2}\right)^x\\ y=C\cdot2^x\] Now, given that \(y(2)=4\), use that information to solve for \(C\).

terenzreignz (terenzreignz):

Cr*p @SithsAndGiggles I thought it was \(\large \ln^2(y)\) and went into a frenzy XD

OpenStudy (anonymous):

@terenzreignz, I thought so too. There's no way to integrate that, I don't think...